KyledaGreat:

Consider the following functions. f(x)=x+53 and g(x)=1x Find the formula for (f∘g)(x) and simplify your answer. Then find the domain for (f∘g)(x). Round your answer to two decimal places, if necessary. (f∘g)(x) =

KyledaGreat:

@vocaloid

Vocaloid:

remember, open circle means take the second function and plug it into the first function so for (f∘g)(x), take g(x) = 1x and plug it into f(x) = x + 53 (the g(x) replaces the x inside f(x))

KyledaGreat:

\[f \frac{ 1 }{ x}\] = (1/x) + 5 / 3

KyledaGreat:

is that right

Vocaloid:

oh, is g(x) = 1/x? if so, yes

KyledaGreat:

for the final answer as it asks (f∘g)(x) = ?

Vocaloid:

yes if g(x) = 1/x, and f(x) = x + 53 then (f∘g)(x) = 1/x + 53 as you determined

KyledaGreat:

it says two decimal places

Vocaloid:

oh, it's also asking for the domain because you have x in the denominator of an expression, x cannot be 0, so it's every real value except 0 idk why they're asking to round to 2 decimal places though, I'm not sure what you're supposed to enter

KyledaGreat:

it's not letting me put 1/x + 53 , it said incomplete

KyledaGreat:

right

Vocaloid:

function seems right, I don't know what their issue is domain is all real values except 0, in set notation it's (-∞,0) ∪ (0,∞)

KyledaGreat:

Given f(x), find g(x) and h(x) such that f(x)=g(h(x)) and neither g(x) nor h(x) is solely x. f(x)=−x^√-x^3−5+4 g(x) = h(x) =

Vocaloid:

possible to get an image of this? having trouble parsing this function

KyledaGreat:

yes, a bit as it is difficult really

KyledaGreat:

i can't get it , i get mixed up

Vocaloid:

can you take a screengrab of the function, or use latex to type it out exactly as it appears?

KyledaGreat:

1 attachment
Vocaloid:

oh this isn't too bad, actually let's let h(x) be the inner part of the root function (-x^3 - 5), that way we can make g(x) simply \[\sqrt[3]{(x)}+4\] notice how when you plug in h(x) = -x^3 - 5 into that g(x), you get g(x) = (cube root of -x^3 - 5) + 4 which is your f(x)

KyledaGreat:

-x^3 + 3sqrtx - 1

Vocaloid:

is this a new problem or are you still attempting the previous one

KyledaGreat:

no , i thought you meant simplify 3sqrt(x) + 4 above and that was somewhat the answer

Vocaloid:

when it asks "Given f(x), find g(x) and h(x) such that f(x)=g(h(x)) " it is asking for two separate functions, g(x) and h(x). once you figure out what your g(x) and h(x) are, that's it, enter what you got as your g(x) and h(x)

KyledaGreat:

f(x)=3-x^3-5+4

Vocaloid:

remember, it is asking for g(x) and h(x) it already tells you what f(x) is

KyledaGreat:

h(x) = -x^3 - 5

Vocaloid:

good, and g(x)?

KyledaGreat:

g(x) = 3sqrtx-5

Vocaloid:

almost notice how f(x) = cube root(-x^3 - 5) + 4 if you choose h(x) = -x^3 - 5 (the expression inside the cube root) then g(x) has to be cube root(x) + 4, not -5, in order to get f(x)

KyledaGreat:

that's right

KyledaGreat:

g(x) = 3sqrt+4

Vocaloid:

g(x) = \[\sqrt[3]{x}+4\] make sure to put the cube root, not square root

KyledaGreat:

yes sir , oh excuse me are you a male or female to say that too ?

KyledaGreat:

if you don't mind me asking

Vocaloid:

I'm a woman

KyledaGreat:

oh , i apologize . forgive me .

KyledaGreat:

Yes ma'am

KyledaGreat:

I tried to put it in the answers but it told me this otherwise

KyledaGreat: